# Field Matters: A lightweight LLM-enhanced Method for CTR Prediction

### v1

@@ 3.2 Feature Representation Enhancement (FRE)
$\textbf{h}^{{}^{\prime}}_{k}=\text{Adaptor}(\textbf{h}_{k}),$ (9)
where $\text{Adaptor}(\cdot)$ can be implemented as a simple linear transformation. We then align feature embeddings with their corresponding field embeddings through normalized KL-divergence [33]:
$\mathbf{h}^{{}^{\prime}}_{k}\leftarrow\text{Softmax}(\mathbf{h}^{{}^{\prime}}_{k}),\ \mathbf{e}_{ik}\leftarrow\text{Softmax}(\mathbf{e}_{ik}),$ (10)
$\mathcal{L}_{KL}=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{D}_{\text{KL}}(\mathbf{e}_{ik},\mathbf{h}^{{}^{\prime}}_{k})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{e}_{ik}\cdot\ln({\mathbf{e}_{ik}}/{\mathbf{h}_{k}^{{}^{\prime}}}).$ (11)
Specifically, here we employ normalized KL-divergence, as it eliminates the influence of embedding magnitude and focuses on aligning the underlying distributions. We also experimented with other commonly used alignment losses, such as contrastive loss, but found that normalized KL-divergence consistently yields superior performance (Appendix B.2). By minimizing $\mathcal{L}_{KL}$, feature embeddings are able to absorb semantic knowledge from the field embeddings, resulting in higher-quality representations. Intuitively, the field semantic embeddings serve as prototypes, encouraging the feature embeddings close to these semantic centers during training.
@@
$\mathcal{L}=\mathcal{L}_{BCE}+\lambda_{kl}\mathcal{L}_{KL},$ (12)
where $\lambda_{kl}$ serves as a hyperparameter balancing the contributions of the two components.
3.3 Feature Interaction Enhancement (FIE)
Field information has been shown beneficial to improve feature interaction modeling in CTR tasks [24, 15, 18]. However, these traditional method only utilize field ID information by employing implicit embedding layers [24] or learnable parameters matrices [15, 18], without taking the textual information of fields into account. Field semantic information provides new insights into feature interactions. For example, semantically, interactions between the features belonging to "user income" and "item price" are likely highly useful for click prediction.
@@ 3.3 Feature Interaction Enhancement (FIE)
$m_{ij}^{F}=\text{FieldInteraction}(\mathbf{h}_{i},\mathbf{h}_{j}),$ (13)
[…]
@@
$\Phi^{new}(\mathbf{X}_{i})=\Phi(\mathbf{X}_{i})+\lambda_{fm}\cdot\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}^{T}\text{>}m_{kl}^{{}^{\prime}}}}})},$ (14)
@@
Notably, our proposed feature interaction enhancement strategy can be seamlessly integrated into a wide range of existing CTR models. This generality motivates our approach of introducing a supplementary field-aware feature interaction component, rather than modifying the inherent feature interaction architecture of each model. In our experiments, we incorporate our LLaCTR module into six representative CTR methods, and observe substantial performance improvements across almost all cases. Detailed results are presented in Section 4, and the implementation details are provided in the Appendix A.4.
4 Experiments
4.1 Experimental Setup
Datasets. We adopt four widely-used public datasets in LLM-enhanced CTR task: Amazon Gift Cards, Amazon Video Games, Amazon Digital Music and MovieLens-1M. We closely refer to [11, 8] and partition each dataset into training, validation, and testing sets with ratios of 8:1:1 according to the timestamp. The details of the above datasets are summarized in Appendix A.1.
Metrics. Following previous work [8, 34, 35], we employed two widely-used metrics LogLoss (binary cross-entropy loss) and AUC (Area under the ROC curve) to evaluate performance. RelaImpr [36], the relative AUC improvement compared to the basic CTR model [28, 7] is also reported. The readers may refer to Appendix A.2 for more details.
Baselines. For comparisons, we selected KAR (RecSys’24) [8], LLM-CF (CIKM’24) [9], CTRL (TORS’23) [7] and EASE (CIKM’24) [10] as LLM-enhanced baselines. These baselines are closely related and representative LLM-enhanced CTR methods. For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B [3] as the LLM. We also include two representative PLM-enhanced methods - CELA (arXiv’24) [12] and ClickPrompt (WWW’24) [11] for comparisons.
We integrated these baseline methods into the following representative traditional CTR models: 1) Classic FM-based methods: FM (ICDM’10) [15] and DeepFM (IJCAI’17) [16]; 2) field-based CTR methods: FwFM (WWW’18) [17] and FmFM (WWW’21) [18]; 3) The state-of-the-art CTR methods: FinalMLP (AAAI’23) [19] and WuKong (ICML’24) [20]. The readers may refer to Appendix A.4 for more details about these methods.
Hyperparameter settings. A grid search is utilized to find the optimal hyperparameters. For all compared methods, we closely refer to the configurations provided in their respective publications to ensure their optimal performance. For LLaCTR, $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}, and the feature sampling number in fine-tuning is set at 1000. The detailed hyperparameter settings are provided in Appendix A.5.
4.2 Experimental Results
In this section, we primarily focus on performance comparison, efficiency analysis, and ablation studies. Additional experimental results, including hyperparameter sensitivity analyses and case studies, are provided in Appendix B.
Table 1: Performance comparisons of LLaCTR with existing PLM/LLM-enhanced methods on six CTR backbone models. The results of PLM-enhanced methods are reported on their best backbones. The best result is bolded and the blue-colored zone indicates that LLaCTR is better than the basic CTR backbone. The line "Average RelaImpr" indicates the average RelaImpr of LLaCTR over six CTR backbones. The mark "*" indicates the improvement is statistically significant ($p$-value $<0.05$).
[…]
@@ 4.2 Experimental Results
[…]
Comparing with traditional CTR models. From the Table 1, we observe LLaCTR can be applied to various types of basic CTR models, and yield performance gains in most cases. The improvements brought by LLaCTR are impressive, achieving an average AUC relative improvement of 2.24% over all basic CTR models across four datasets. These results demonstrate the effectiveness of LLaCTR by injecting field semantic knowledge.
Comparing with PLM/LLM-enhanced CTR models. We observe that LLaCTR outperformed two representative PLM-enhanced CTR models (CELA and ClickPrompt) and four LLM-enhanced CTR models (KAR, LLM-CF, CTRL and EASE) in most cases (89%), which validates the effectiveness of our field-level enhancement paradigm. This result is highly surprising, as we just use the lightweight field knowledge rather than the cumbersome instance-level (or user/item-level) knowledge utilized by recent work. The reason may be : 1) field knowledge is indeed important and helpful for CTR prediction; 2) the lightweight field knowledge has been effectively exploited, enhancing both feature representation and feature interactions.
Efficiency Analysis. The AUC performance and the total training time of compared methods are shown in Figure 4. The performance of LLM-enhanced methods are reported on their best backbones. We observe: 1) Compared to three representative LLM-enhanced methods, LLaCTR can improve efficiency by 10-100x and achieve better performance. 2) With the increase of sampling number in the SSFT module, both the training time and AUC of LLaCTR will increase. This indicates increasing the number of samples used for self-supervised learning can yield higher-quality field semantic knowledge, but require more training time cost. 3) Even without fine-tuning, LLaCTR still achieves decent results. This suggests our LLaCTR could be applied in resource-constrained scenarios, enhancing existing CTR methods with incurring limited additional time cost.
Ablation Study. We further conduct the ablation study, where the Self-supervised Fine-tuning (SSFT) module, Feature Representation Enhancement (FRE) module or Feature Interaction Enhancement (FIE) module has been removed respectively. The results are presented in Table 2. As can be seen, both the three components are important — with few exceptions, removing SSFT (w/o FT), FRE (w/o FRE) or FIE ($\lambda_{fm}$) module would result in performance drops.
@@
Figure 4: AUC and training time of compared methods. The “sam-num” represents the sampling feature number for each field of LLaCTR, and "w/o FT" represents fine-tuning has been removed.
Table 2: Ablation study on LLaCTR with DeepFM and WuKong backbones. The results on other backbones are presented in Appendix B.3.
[…]
w/o FT 0.4687 0.6759 0.5771 0.6732 0.3393 0.7683 0.4064 0.8352
[…]
5 Related Work
Traditional CTR Prediction. CTR Prediction serves as a core function module in personalized online services [37, 1, 38, 39, 16, 40]. The key idea of traditional CTR prediction models is to capture the feature interaction patterns, which indicates the combination relationships of multiple features. The feature interaction layer is generally implemented by a combination of various operations. Product operators include classical bi-level models [15, 41, 42] and higher-order [43, 44, 45, 46, 47] models. Convolutional operators [48, 27, 26] utilize CNN [49] and GCN [50] to capture the implicit high order interaction. Due to the success of attention mechanism, attention-based methods [51, 21, 34, 35, 28, 29] also achieved great success. In recent years, several field-aware CTR methods have been proposed [24, 17, 18], however, these methods relied solely on field IDs as categorical indicators without leveraging the rich semantic information inherent in field descriptions.
LLMs as CTR predictors. Initial efforts explored the usage of pre-trained language models as CTR predictors by reformulating recommendation tasks as natural language [4, 52, 53]. With the advent of LLMs and their remarkable abilities in content comprehension and semantic reasoning [2, 3], recent efforts have shifted to explore using larger-scale Language Models to generate CTR prediction results [54, 55, 56, 57, 58, 59, 60, 31], e.g., TALLRec [5] uses instruction tuning with user’s historical interactions, enabling LLMs to achieve decent CTR prediction capability in zero-shot and few-shot scenarios, while HLLM [61] utilizes two independent LLMs to model users and items separately, and jointly trains them with both generative and discriminative recommendation losses. However, the high inference latency and resource demands of LLMs poses significant obstacles to their deployment in real-world large-scale recommender systems. Although recent studies have proposed various acceleration strategies, e.g., BAHE [6] freezes the shallow layers of LLMs to encode users’ atomic interactions and pre-stores them offline to reduce the number of parameters involved in training and inference, while Rella [62] reduces inference overhead by shortening user interaction sequences through the elimination of items that are semantically irrelevant to the target item, the latency and resource consumption nonetheless remain substantial due to the LLMs’ large parameter counts and complex architectures.
LLM-enhanced CTR Prediction. To address the efficiency challenges inherent in LLMs-based CTR predictors, the LLM-enhanced CTR paradigm has been extensively investigated as an alternative approach. This paradigm primarily focus on leveraging the powerful semantic encoding and knowledge reasoning capabilities of LLMs to enhance traditional CTR models [63, 8, 9, 7, 10]. For example, KAR [8] leverages the LLM as a open-word knowledge base to enhance the profile of users and items. LLM-CF [9] uses LLMs to generate a chains of thought (CoT) for sampled instances, which is further retrieved to enhance CTR models. CTRL [7] utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align the knowledge. EASE [10] trains a semantic adaptor to align item embeddings with LLM, and use it for encoding during inference. Other work also focuses on specific CTR scenarios, such as cross-domain[64, 65] and interpretability [66, 67]. Despite their promising performance, they involve enhancements at the instance-level or user/item-level, which results in great inefficiency issue in dealing with large-scale instances or entities in practice. Our approach is applied on the more efficient and cost-effective field-level, without incurring superabundant computational overhead.
There are also two related topics: 1) PLM-enhanced CTR methods: Early studies have leveraged pretrained language models (PLMs, like BERT [68] and RoBERTa[69]) to enhance CTR models via encoding textual features [70, 71, 72, 73, 53, 11, 12], e.g., CELA [12] pretrains a PLM on item descriptions and aligns item textual embeddings with ID-based embeddings for enhancement. ClickPrompt [11] uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge. However, the capacity of early PLMs is generally inferior to that of recent LLMs, and recent studies have demonstrated the superior performance of LLM-enhanced CTR methods. Consequently, our work aims to further investigate and advance the state-of-the-art in LLM-enhanced CTR modeling. 2) LLM-enhanced recommendation: Recent years also witnessed the integration of LLMs on other recommendation tasks,e.g., sequential recommendation [74, 75, 59, 76, 77, 78], collaborative filtering [79, 80, 65, 81], cross-domain recommendation [82, 83]. It is important to note that CTR prediction differs from these tasks in that it typically involves various features and emphasizes modeling feature interactions. Thus, our LLaCTR is specifically designed to leverage LLMs for improved feature representation and interactions in CTR prediction.
6 Conclusion and Limitations
This work proposes a novel lightweight method, LLaCTR, which leverages LLMs to extract semantic knowledge from feature fields for enhancing CTR models. Specifically, LLaCTR employs a self-supervised field-feature fine-tuning to distill high-quality field-level knowledge, which is subsequently utilized to improve both feature representation and feature interaction modeling. Extensive experiments demonstrate the superiority of LLaCTR in both predictive accuracy and computational efficiency. Nevertheless, a current limitation of LLaCTR is that it leverages field-level semantic knowledge to guide bi-level feature interaction modeling. Exploring more advanced strategies for incorporating field knowledge to capture higher-order feature interactions remains an interesting and promising direction for future research.
Appendix A Experimental Details
A.1 Datasets
We conduct experiments on four real-world datasets from different recommendation scenarios. The datasets statistics are summarized in Table A.1.
Table A.1: Statistics of the datasets.
[…]
#Feature 1,981,330 60,119,995 1,695,642 9,001,881
#User 132,732 2,766,656 100,952 6,040
#Item 1,137 137,249 70,511 3,706
#Interaction 152,410 4,624,615 130,434 1,000,209
Density $1.01\times 10^{-3}$ $1.22\times 10^{-5}$ $1.83\times 10^{-5}$ $4.47\times 10^{-2}$bigstrut[b]
@@ A.1 Datasets
  • Amazon Review Datasets [^1] https://amazon-reviews-2023.github.io/ are widely-used e-commercial datasets [9, 7, 11, 84] from Amazon website with ratings ranging from 1 to 5. We use the Gift Cards, Video Games and Digital Music subsets in our experiments. Following the standard practice in Lin et al. [11], we binarize the ratings with a threshold of 4 and use 5-core setting, i.e., all users and items have at least 5 interactions.
@@
  • MovieLens-1M Dataset [^2] https://grouplens.org/datasets/MovieLens/1m/ is a movie recommendation dataset from Movielens website with ratings ranging from 1 to 5. We binarize the ratings with a threshold of 4, while removing neutral samples with ratings equal to 3 following previous work [7].
@@
Following [11], we organized the review behaviors in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1 after preprocessing. All the experiments are implemented with PyTorch and run on 4x Nvidia A800 GPUs.
A.2 Metrics
To evaluate the performance of CTR prediction methods, we closely follow [28, 8, 11, 7] and adopt AUC (area under the ROC curve), LogLoss (binary cross-entropy loss) and RelaImpr (AUC relative improvement) as the evaluation metrics:
  • AUC [85]: AUC is the area under the ROC curve and measures the goodness of order by ranking all the items with predicted CTR. For a random guesser, the value of AUC is 0.5.
  • LogLoss: Logloss is the logarithmic loss value of binary cross-entropy loss and quantifies the accuracy of predicted probabilities by penalizing false classifications.
  • RelaImpr [36]: RelaImpr measures the AUC relative improvement over basic CTR models and is defined as below:
    $RelaImpr=(\frac{\text{AUC (measured\ model)}-0.5}{\text{AUC (base\ model)}-0.5}-1)\times 100\%.$ (15)
Slightly higher AUC or lower Log Loss (e.g., 0.001) can be regarded as significant improvement in CTR prediction, as indicated by previous studies [34, 35].
A.3 Baselines

### v2

@@ 3.2. Feature Representation Enhancement (FRE)
(9) $\textbf{h}^{{}^{\prime}}_{k}=\text{Adaptor}(\textbf{h}_{k}),$
where $\text{Adaptor}(\cdot)$ can be implemented as a simple linear transformation. We then align feature embeddings with their corresponding field embeddings through normalized KL-divergence (Tian et al., 2019):
(10) $\mathbf{h}^{{}^{\prime}}_{k}\leftarrow\text{Softmax}(\mathbf{h}^{{}^{\prime}}_{k}),\ \mathbf{e}_{ik}\leftarrow\text{Softmax}(\mathbf{e}_{ik}),$
(11) $\small\mathcal{L}_{KL}=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{D}_{\text{KL}}(\mathbf{e}_{ik},\mathbf{h}^{{}^{\prime}}_{k})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{e}_{ik}\cdot\ln({\mathbf{e}_{ik}}/{\mathbf{h}_{k}^{{}^{\prime}}}).$
Specifically, here we employ normalized KL-divergence, as it eliminates the influence of embedding magnitude and focuses on aligning the underlying distributions. We also experimented with other commonly used alignment losses, such as contrastive loss, but found that normalized KL-divergence consistently yields superior performance (see Appendix B.2.2 for details). By minimizing $\mathcal{L}_{KL}$, feature embeddings are able to absorb semantic knowledge from the field embeddings, resulting in higher-quality representations. Intuitively, the field semantic embeddings serve as prototypes, encouraging the feature embeddings close to these semantic centers during training.
@@
(12) $\mathcal{L}=\mathcal{L}_{BCE}+\lambda_{kl}\mathcal{L}_{KL},$
where hyperparameter $\lambda_{kl}$ balances their contributions.
3.3. Feature Interaction Enhancement (FIE)
Field information has been shown beneficial to improve feature interaction modeling in CTR tasks (Juan et al., 2016; Rendle, 2010; Sun et al., 2021). However, these traditional method only utilize field ID information by employing implicit embedding layers (Juan et al., 2016) or learnable parameters matrices (Rendle, 2010; Sun et al., 2021), without taking the textual information of fields into account. Field semantic information provides new insights into feature interactions. For example, semantically, interactions between the features belonging to "user income" and "item price" are likely highly useful for click prediction.
@@ 3.3. Feature Interaction Enhancement (FIE)
(13) $m_{ij}^{F}=\text{FieldInteraction}(\mathbf{h}_{i},\mathbf{h}_{j}),$
where $\text{FieldInteraction}(\cdot)$ denotes the field interaction layer, which can be implemented by various architecture. In practice, we find that cosine similarity is effective — intuitively, the fields with similar semantic may provide more insights on prediction. For convenience, we collect all pairwise field interaction scores into a matrix $\mathbf{M}^{F}$. Subsequently, we employ a learnable adaptor to re-scale these importance scores, i.e., $\mathbf{M}^{{}^{\prime}}=\text{Adaptor}(\mathbf{M}^{F})$, where the adaptor can be simply implemented via a linear layer.
@@
(14) $\small\Phi^{new}(\mathbf{X}_{i})=\Phi(\mathbf{X}_{i})+\lambda_{fm}\cdot\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}^{T}\text{>}m_{kl}^{{}^{\prime}}}}})},$
@@
Notably, our proposed feature interaction enhancement strategy can be seamlessly integrated into a wide range of existing CTR models. This generality motivates our approach of introducing a supplementary field-aware feature interaction component, rather than modifying the inherent feature interaction architecture of each model. In our experiments, we incorporate our LLaCTR module into six representative CTR methods, and observe substantial performance improvements across almost all cases. Detailed results are presented in Section 4, and the implementation details are provided in the Appendix A.
Table 1. Statistics of the datasets.
  Dataset Gift Cards Video Games Digital Music MovieLens-1M
#Field 13 13 13 8
#Feature 1,981,330 60,119,995 1,695,642 9,001,881
#User 132,732 2,766,656 100,952 6,040
#Item 1,137 137,249 70,511 3,706
#Interaction 152,410 4,624,615 130,434 1,000,209
Density $1.01\times 10^{-3}$ $1.22\times 10^{-5}$ $1.83\times 10^{-5}$ $4.47\times 10^{-2}$
Table 2. Performance comparisons of LLaCTR with existing PLM/LLM-enhanced methods. The results of PLM-enhanced methods are reported on their best backbones. The best result is bolded and the blue-colored zone indicates that LLaCTR is better than the basic CTR backbone. The mark "*" indicates the improvement is statistically significant ($p$-value <0.05<0.05).
  Gift Cards Video Games Digital Music       MovieLens-1M
Method Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr
CELA 0.4658 0.6746 1.02% 0.5798 0.6775 0.44% 0.3364 0.7638 0.37% 0.4031 0.8373 0.24%
ClickPrompt 0.4656 0.6754 1.48% 0.5747 0.6799 1.79% 0.3356 0.7652 0.90% 0.4025 0.8379 0.42%
Base 0.4606 0.6621 0.5805 0.6673 0.3359 0.7612 0.4081 0.8354
@@
WuKong LLaCTR 0.4745 0.6802 4.24%* 0.5763 0.6819 2.91%* 0.3377 0.7564 1.99%* 0.4027 0.8397 2.21%*
Average RelaImpr 3.41% 3.65% 1.16% 0.74%
4. Experiments
We aim to answer the following research questions:
  • $\mathbf{RQ1:}$ How does LLaCTR perform compared with the state-of-the-art CTR methods?
  • $\mathbf{RQ2:}$ How is the efficiency of LLaCTR compared with existing LLM-enhanced CTR methods?
  • $\mathbf{RQ3:}$ What are the impacts of different components of LLaCTR?
4.1. Experimental Setup
@@
Figure 5. AUC and training time of compared methods. The “sam-num” is the sampling feature number of each field, and "w/o FT" represents fine-tuning has been removed.
4.1.1. Datasets
Following previous work (Lin et al., 2024a; Xi et al., 2024), we conduct experiments on four widely used public datasets in LLM-enhanced CTR task: Amazon Gift Cards, Amazon Video Games, Amazon Digital Music and MovieLens-1M.
@@ 4.1.1. Datasets
  • Amazon Review Datasets [^1] https://amazon-reviews-2023.github.io/ are well-known e-commercial datasets (Sun et al., 2024; Li et al., 2023; Lin et al., 2024a; Bao et al., 2023a) with ratings ranging from 1 to 5. Following (Lin et al., 2024a), we binarize the ratings with a threshold of 4 and use the 5-core setting, i.e., all users and items have at least 5 interactions.
@@
  • MovieLens-1M Dataset [^2] https://grouplens.org/datasets/MovieLens/1m/ is a benchmark movie recommendation dataset from Movielens with ratings ranging from 1 to 5. We binarize the ratings with a threshold of 4, while removing neutral samples with ratings equal to 3 following (Li et al., 2023).
@@
The dataset statistics are summarized in Table 1. Following (Li et al., 2023; Lin et al., 2024a), we organize the review behaviors in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1 after preprocessing.
4.1.2. Evaluation Metrics
Following previous work (Lin et al., 2024a; Song et al., 2019), we employ two widely-used metrics LogLoss (binary cross-entropy loss) and AUC (area under the ROC curve) to evaluate performance. The RelaImpr (Yan et al., 2014) (relative AUC improvement) is also reported. Notably, slightly higher AUC or lower LogLoss (e.g., 0.1%) can be regarded as significant improvement in CTR prediction, as indicated by previous studies (Huang et al., 2019; Song et al., 2019).
4.1.3. Baselines
For comparisons, we selected KAR (RecSys’24) (Xi et al., 2024), LLM-CF (CIKM’24) (Sun et al., 2024), CTRL (TORS’23) (Li et al., 2023) and EASE (CIKM’24) (Qiu et al., 2024) as LLM-enhanced baselines. These baselines are closely related and representative LLM-enhanced CTR methods. We also include two representative PLM-enhanced methods — CELA (arXiv’24) (Wang et al., 2024) and ClickPrompt (WWW’24) (Lin et al., 2024a) for comparisons.
We integrated these baseline methods into the following representative traditional CTR models: 1) Classic FM-based methods: FM (ICDM’10) (Rendle, 2010) and DeepFM (IJCAI’17) (Guo et al., 2017); 2) Field-based CTR methods: FwFM (WWW’18) (Pan et al., 2018) and FmFM (WWW’21) (Sun et al., 2021); 3) The state-of-the-art CTR methods: FinalMLP (AAAI’23) (Mao et al., 2023b) and WuKong (ICML’24) (Zhang et al., 2024). The readers may refer to Appendix A for more details about these methods.
4.1.4. Implementation Details
We use the Adam (Kingma and Ba, 2014) optimizer to train all the CTR models. The learning rate (lr) is set as 0.001 and the weight decay (wd) is tuned in {1e-2, 1e-3, 1e-4, 1e-5, 1e-6, 0}. We fix the feature embedding size to 32 and batch size to 4096 for all the backbones and datasets. For all compared methods, we closely refer to the configurations provided in their respective publications to ensure their optimal performance. For LLaCTR fine-tuning, we fix the batch size as 128, the learning rate of LLM as 1e-4 and rank of LoRA (Hu et al., 2021) as 8, respectively. During LLaCTR enhancement, the $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}. We set the feature sampling number as 1000 by default, and set temperature parameter $\tau$ of contrastive loss into 0.02 following previous work (Wang et al., 2023). All methods are implemented in PyTorch and run on 8 Nvidia A800 GPUs. More implementation details are provided in the Appendix A.
4.2. Experimental Results
In this section, we focus on performance comparison, efficiency analysis, and ablation studies. Additional results, including time complexity analysis, hyperparameter sensitivity analysis and more details, are provided in Appendix B.
4.2.1. Performance Comparison (RQ1)
Table 2 shows the performance comparison of the proposed LLaCTR against the baseline methods. We observe that:
Comparing with traditional CTR models. LLaCTR can be applied to various types of basic CTR models, and yield performance gains in most cases. The improvements brought by LLaCTR are impressive, achieving an average AUC relative improvement of 2.24% over all basic CTR models across four datasets. These results demonstrate the effectiveness of LLaCTR by injecting field semantic knowledge.
Comparing with PLM/LLM-enhanced CTR models. We observe that LLaCTR outperforms two representative PLM-enhanced CTR models (CELA and ClickPrompt) and four LLM-enhanced CTR models (KAR, LLM-CF, CTRL and EASE) in most cases (89%), which validates the effectiveness of our field-level enhancement paradigm. This result is highly surprising, as we only utilize the lightweight field knowledge rather than the cumbersome instance-level (or user/item-level) knowledge utilized by recent work. The reasons may be 1) field knowledge is indeed important and helpful for CTR prediction; 2) the lightweight field knowledge has been effectively exploited, enhancing both feature representation and feature interactions.
4.2.2. Efficiency Analysis (RQ2)
In this section, we conduct an in-depth analysis on the training efficiency of different LLM-enhanced CTR methods. In terms of inference efficiency, since all these methods pre-store LLM knowledge and ultimately depend on CTR models for recommendations, the variation is much smaller than that observed during training (cf. Table 8 in Appendix B.3).
The AUC performance and the total training time of compared methods are shown in Figure 5. The performances of LLM-enhanced methods are reported on their best backbones. We observe: 1) Compared to three representative LLM-enhanced methods, LLaCTR can improve efficiency by 10-100x and achieve better performance. 2) With the increase of the sampling number in the SSFT module, both the training time and AUC of LLaCTR will increase. This indicates increasing the number of samples used for self-supervised learning can yield higher-quality field semantic knowledge, but require more training time. 3) Even without fine-tuning, LLaCTR still achieves decent results. This suggests our LLaCTR could be applied in resource-constrained scenarios, enhancing existing CTR methods with incurring limited additional time cost. Besides, we also study the efficiency bottleneck and the time complexity of these LLM-enhanced methods, the readers may refer to Appendix B.1 for more details.
4.2.3. Ablation Study (RQ3)
We further conduct the ablation study, where the Self-supervised Fine-tuning (SSFT) module, Feature Representation Enhancement (FRE) module or Feature Interaction Enhancement (FIE) module is removed respectively. The results are presented in Table 3. As can be seen, both the three components are important — removing SSFT, FRE or the FIE module would result in performance drops in most case. Delving deeper into the SSFT module, we observe that self-supervised field-feature fine-tuning on LLMs is indeed helpful. The field description encoded by the fine-tuned LLM has higher quality than that extracted from the pre-trained corpus. For FRE module, we observe that the developed KL-Divergence loss is also important. It can effectively inject semantic knowledge of fields into the feature representations. For FIE module, we find that for both explicit feature modeling and implicit feature modeling CTR methods, injecting field interaction scores can effectively guide feature interaction and bring AUC improvement.
Table 3. Ablation study on DeepFM and WuKong. Results on the other backbones are shown in Appendix B.4.
                         Method Gift Games Music Movie
DeepFM Base 0.6713 0.6692 0.7582 0.8342
w/o FT 0.6759 0.6732 0.7683 0.8352
w/o $\lambda_{kl}$ 0.6764 0.6699 0.7582 0.8343
w/o $\lambda_{fm}$ 0.6787 0.6741 0.7685 0.8351
LLaCTR 0.6797 0.6744 0.7687 0.8354
WuKong Base 0.6728 0.6767 0.7514 0.8323
w/o FT 0.6800 0.6777 0.7558 0.8322
w/o $\lambda_{kl}$ 0.6730 0.6784 0.7524 0.8332
w/o $\lambda_{fm}$ 0.6763 0.6783 0.7494 0.8347
LLaCTR 0.6802 0.6819 0.7564 0.8397
5. Related Work
5.1. Traditional CTR Prediction
CTR prediction is a core functional module in personalized online services, where the key idea is to capture feature interaction patterns that capture the combinational relationships among multiple features. Traditional methods employ various operations for feature interaction, including product-based operators (Cheng et al., 2016; Wang et al., 2021), convolutional networks (Liu et al., 2015), and attention mechanisms (Huang et al., 2019). Recently, field-aware CTR methods (Pan et al., 2018; Sun et al., 2021) have gained attention, though they primarily rely on field IDs as categorical indicators without fully leveraging the semantic richness of field descriptions.
5.2. LLMs as CTR predictors
Initial efforts explored the pre-trained language models as CTR predictors by reformulating recommendation tasks as NLP problems (Geng et al., 2022; Cui et al., 2022; Liu et al., 2022). With the advent of LLMs in content comprehension and semantic reasoning (Achiam et al., 2023; Dubey et al., 2024), LLMs have demonstrated strong potential in CTR prediction (Bao et al., 2023b; Chen et al., 2024). However, LLMs suffer from high inference latency and resource demands. While acceleration techniques exist (e.g., BAHE (Geng et al., 2024) for parameter reduction, Rella (Lin et al., 2024b) for sequence shortening), their computational overhead remains prohibitive for real-world deployment due to massive model sizes.
5.3. LLM-enhanced CTR Prediction
To address the efficiency challenges inherent in LLM-based CTR predictors, the LLM-enhanced CTR paradigm has been extensively investigated as an alternative approach. This paradigm primarily focuses on leveraging the powerful semantic encoding and knowledge reasoning capabilities of LLMs to enhance traditional CTR models (Li et al., 2023; Qiu et al., 2024). For example, KAR (Xi et al., 2024) leverages LLMs to enhance the knowledge of users and items. LLM-CF (Sun et al., 2024) uses LLMs to generate chains of thought (CoT) for enhancement. CTRL (Li et al., 2023) utilizes the LLM as the feature encoder for instances and uses contrastive learning to align the knowledge. EASE (Qiu et al., 2024) trains a semantic adaptor to align item embeddings with LLM. Other work also focuses on specific CTR scenarios, such as cross-domain(Fu et al., 2023) and interpretability (Yu et al., 2024). Despite their promising performance, they involve enhancements at the instance-level or user/item-level, which results in great inefficiency issue in dealing with large-scale instances or entities in practice. Our approach is applied on the more efficient and cost-effective field-level, without incurring superabundant computational overhead.
5.4. Other Related Work
There are also two related topics: 1) PLM-enhanced CTR methods: Early studies have leveraged pretrained language models (PLMs, such as BERT (Devlin et al., 2019)) to enhance CTR models, e.g., CELA (Wang et al., 2024) pretrains PLMs on item features and aligns them with ID embeddings for enhancement. ClickPrompt (Lin et al., 2024a) uses the ID embeddings as prefix soft tokens in PLMs for generating instance-level knowledge. However, the capacity of early PLMs is generally inferior to that of recent LLMs, and recent studies have demonstrated the superior performance of LLM-enhanced methods. Consequently, our work aims to further investigate and advance the LLM-enhanced CTR modeling. 2) LLM-enhanced recommendation: Recent years have also witnessed the integration of LLMs on other recommendation tasks (Cui et al., 2025; Wang et al., 2025a, 2026), e.g., sequential recommendation (Cui et al., 2024; Wang et al., 2025b), collaborative filtering (Ren et al., 2024; Lin et al., 2025). Notably, CTR prediction differs from these tasks in that it typically involves various features and emphasizes modeling feature interactions. Thus, our LLaCTR is specifically designed to improve feature representation and interactions in CTR prediction.
6. Conclusion
This work proposes a lightweight method, LLaCTR, which leverages LLMs to extract semantic knowledge from feature fields for enhancing CTR models. Specifically, LLaCTR employs a self-supervised field-feature fine-tuning to distill high-quality field-level knowledge, which is subsequently utilized to improve both feature representation and feature interaction modeling. Extensive experiments demonstrate the superiority of LLaCTR in both predictive accuracy and computational efficiency. In the future, it would be valuable to investigate more sophisticated strategies for integrating field semantics to further enhance CTR methods.
Acknowledgements.
This work is supported by the Zhejiang Province “JianBingLingYan+X” Research and Development Plan (2025C02020). We thank the reviewers for their valuable and insightful suggestions that improve the paper.
Appendix A Appendix A: Experimental Details
A.1. Baselines